\begin{thebibliography}{100}

\bibitem{Sciama1} D. W. Sciama, P. C. Waylen and R. C. Gilman,
{\sl Phys. Rev.} {\bf 187} (1969) 1762.
\bibitem{Thorne} K. S. Thorne, {\sl Rev. Mod. Phys.} {\bf 52} (1980) 299.
\bibitem{Kramer} D. Kramer, H. Stephani, E. Herlt and 
M. MacCallum, {\sl Exact Solutions of Einstein's Field Equations},
(1980) Cambridge University Press, Cambridge.
\bibitem{Palle1} D. Palle, {\bf arXiv:astro-ph/0211481} (2002).
\bibitem{Cooperstock} F. I. Cooperstock, {\sl Mod. Phys. Lett.}
{\bf A 14} (1999) 1531; {\sl Ann. Phys. (N.Y.)} {\bf 282} (2000) 115.
\bibitem{Dirac} P. A. M. Dirac, {\sl Proc. Roy. Soc. (London)}
{\bf A 126} (1930) 360.
\bibitem{Englert} F. Englert and R. Brout, {\sl Phys. Rev. Lett.}
{\bf 13} (1964) 321.
\bibitem{Higgs} P. W. Higgs, {\sl Phys. Rev. Lett.} {\bf 13} (1964) 508.
\bibitem{BPST} A. A. Belavin, A. M. Polyakov, A. S. Schwartz and
Yu. S. Tyupkin, {\sl Phys. Lett.} {\bf 59 B} (1975) 85.
\bibitem{tHooft} G. 't Hooft, {\sl Phys. Rev.} {\bf D 14} (1976) 3432.
\bibitem{PQ} R. D. Peccei and H. R. Quinn, {\sl Phys. Rev.} {\bf D 16}            (1977) 1791.
\bibitem{Weinberg1} S. Weinberg, {\sl Phys. Rev. Lett.} {\bf 40} (1978) 223.
\bibitem{Palle2}
  D. Palle,
  {\sl Nuovo Cimento} {\bf A 109} (1996) 1535.
\bibitem{Mack} H. A. Kastrup, {\sl Ann. Phys. (Leipzig)} {\bf 9} (1962)
 388; G. Mack and A. Salam, {\sl Ann. Phys. (N.Y.)} {\bf 53} (1969) 174.
\bibitem{Palle3} D. Palle, {\bf arXiv:hep-ph/0703203} (2007).
\bibitem{Palle4}
  D. Palle,
  {\sl Hadronic Journal} {\bf 24} (2001) 87 (2001);
  ibidem {\bf 24} (2001) 469.
\bibitem{Sciama2} D. W. Sciama, {\sl On the analogy between charge
and spin in general relativity} (1962) in {\sl Recent Developments
in General Relativity}, Pergamon Press, Oxford, p.415:
T. W. B. Kibble, {\sl J. Math. Phys.} {\bf 2} (1961) 212.
\bibitem{Weinberg2}
  S. Weinberg,
  {\sl Gravitation and Cosmology} (1972)
  J. Wiley and Sons, New York.
\bibitem{Trautman}
  A. Trautman,
  {\sl Nature} {\bf 242} (1973) 7.
\bibitem{Palle5}
  D. Palle,
  {\sl Nuovo Cimento} {\bf B 111} (1996) 671.
\bibitem{Hawking}
   S. W. Hawking, {\sl Commun. math. Phys.} {\bf 43} (1975) 199.
\bibitem{Weinberg3} S. Weinberg, {\sl Rev. Mod. Phys.} {\bf 61} 
   (1989) 1.
\bibitem{Palle6}
  D. Palle,
  {\sl Nuovo Cimento} {\bf B 114} (1999) 853;
  ibidem {\bf B 121} (2006);
  ibidem {\bf B 122} (2007) 67.
\bibitem{Palle7}
  D. Palle, {\bf arXiv:0802.2060} (2008).
\bibitem{Mach}
  E. Mach, {\sl Die Mechanik in ihrer Entwicklung historisch-kritisch
  dargestellt} (1883) Brockhaus, Leipzig.
\bibitem{Palle8}
  D. Palle,
  {\sl Nuovo Cimento} {\bf B 115} (2000) 445.
\bibitem{Basilakos}
 S. Basilakos, J. C. Bueno Sanchez and L. Perivolaropoulos,
  {\bf arXiv:0908.1333} (2009).
\bibitem{Bean}
 R. Bean, {\bf arXiv:0909.3853} (2009).
\bibitem{Palle9}
  D. Palle, {\bf arXiv:0902.1852} (2009).
\bibitem{Tevatron}
  The D0 Collaboration, {\sl Phys. Rev.} {\bf D 64} (2001) 032003.
\bibitem{Palle10}
  D. Palle, 
{\sl Nuovo Cimento} {\bf B 118} (2003) 747.
\bibitem{Palle11}
  D. Palle, 
{\sl Nuovo Cimento} {\bf B 119} (2004) 1121.
\bibitem{Birch}
  P. Birch,
{\sl Nature} {\bf 298} (1982) 451; ibidem {\bf 301} (1982) 736.
\bibitem{Longo}
M. J. Longo, {\bf arXiv:0908.1732} (2009).
\bibitem{Kashlinsky}
A. Kashlinsky, F. Atrio-Barandela, D. Kocevski and H. Ebeling,
  {\sl Astrophys. J. Lett.} {\bf 686} (2008) L49;
 R. Watkins, H. A. Feldman and M. J. Hudson,
  {\sl M.N.R.A.S.} {\bf 392} (2009) 743;
 G. Lavaux, R. B. Tully, R. Mohayaee and
   S. Colombi,
   {\bf arXiv:0810.3658} (2008);
A. Kashlinsky, F. Atrio-Barandela, H. Ebeling, A. Edge and D. Kocevski,
   {\bf arXiv:0910.4958} (2009).
\bibitem{Palle12}
  D. Palle,
 {\sl Acta Physica Slovaca} {\bf 55} (2005) 237.
\bibitem{Palle13}
  D. Palle,
 {\sl Acta Physica Slovaca} {\bf 55} (2005) 549.
\bibitem{Palle14}
  D. Palle,
{\sl Nuovo Cimento} {\bf B 112} (1997) 943.

\end{thebibliography}

\documentclass[aps,prd,preprint,a4paper]{revtex4} 
\begin{document}
\title{Comments on the theory of noncontractible space and the misuse of
       the theory of vacuum}
\author{Davor Palle}
\affiliation{
Zavod za teorijsku fiziku Instituta Rugjer Bo\v skovi\' c \\
Bijeni\v cka cesta 54, HR-10002 Zagreb, Croatia; e-mail:palle@irb.hr}

\date{November 11, 2009}

\begin{abstract}
I comment critically on the use and misuse of the theory of vacuum,
pseudoparticles and pseudotensors. The mathematical and phenomenological
arguments against the Higgs mechanism and the inflationary scenario
are presented. I conclude with a proposal on how to mathematically 
improve our understanding of fundamental interactions, basing it on the
theory of noncontractible space. The nature of the force of gravity 
appears crucial in my reasoning.
\end{abstract}

\maketitle 

\section{Introduction}

In this comment, I should like to address a few important issues
in field theory with its implications in particle physics and
cosmology. Understanding the dynamics of the largest and the smallest 
entities of the Universe requires a formulation of theories that
can explain the appearance of physical structures, physical forces 
and all the observed phenomena. One can say nowadays that cosmology  
represents literally the theory of everything, with measurements and
observations made in the laboratory or at the outskirts of the Universe.

The natural question arises whether one can be satisfied 
with the present, widely accepted 
theoretical mechanisms that attempt to resolve the
fundamental questions in physics.

My answer to this question is firmly negative. Why?
Let me explain it in the next section.

\section{Dissatisfaction with standard models}

Despite the fact that we can describe the world of elementary particles
by the non-Abelian gauge theory $SU(3)\times SU(2)_{L}\times U(1)$,
there is a lot of unanswered questions. Particles and interactions
are described by local quantum fields in the form of the second quantization
in Minkowski spacetime of the special theory of relativity.
The two basic principles, relativity and quantization, applied
to the perturbation theory of non-Abelian gauge interactions 
allow to confront theory with experiments,giving 
us the confidence in the standard model $SU(3)\times SU(2)_{L}\times U(1)$.
Calculations of the QCD bound state spectra with lattice gauge theory 
improve further our understanding of the world of elementary particles.

However, there are plenty of reasons to be dissatisfied with
the standard model: the problem of ultraviolet infinities is
resolved only in perturbative approach by the theory of
renormalization: neutrinos are massless Dirac particles,
the scalar Higgs sector is a theoretical construct with arbitrarily
inserted parameters;
all masses are free parameters; there is no candidate for a
dark matter particle; no baryon- or lepton-number violating
interactions that are required by cosmological considerations;
no link whatsoever with the theory of gravity.

The standard model in cosmology is called concordance model, with
the following ingredients: Einstein's theory of general relativity
can describe the Universe expanding from the initial singularity in
the form of Friedmann-Robertson-Walker geometry; the flatness and
horizon problems, as well as the structure formation problem, are 
claimed to be 
resolved by adding an inflaton scalar field whose dynamics is defined
by various types of potentials. 

The idea to introduce ad hoc in cosmology a scalar field instead of  
a certain quantum theory
of gravity at distances close to Planck scale, is very pragmatic and
unusual. However, the consequences are far-reaching:
similarly as in the scalar Higgs sector in particle physics,
all parameters of any version of inflationary cosmology are
free unknown parameters; inflation cannot predict mass density or
the cosmological constant but only total mass density of the
Universe; primordial spectrum of density contrast is again given
by unknown potential of the inflaton scalar; as a consequence of
the Planck and GUT scales dynamics of the inflaton, there is an imprint
of gravitational waves in the tensor mode of the fluctuations
of the CMB temperature field. The cosmological principle also assumes 
perfect homogeneity and perfect isotropy built in 
Friedmann-Robertson-Walker geometry, because there are
no mechanisms to break them in a natural way.

That was just a brief overview of the problems with which standard
models in particle physics and cosmology are confronted.

\section{Unification programs and gravity}

There is a general agreement that the force of gravity is also
a local gauge force in analogy with the $SU(3)\times SU(2)_{L}\times U(1)$ unitary
groups gauge forces. Common wisdom asserts at the same time that general relativity
is incompatible with quantum mechanics.
Let us analyze these statements in more detail.

Einstein's theory of gravity or general relativity (GR) is singular at
vanishing distances when collapse of astrophysical objects occurs or 
at the Big Bang cosmological initial singularity. 
The same singularity occurs in the nonrelativistic Newton's theory of gravity. Maxwell's theory of classical electrodynamics,
as a classical gauge field theory of electromagnetism, is singular
because if one attempts to compute self-energy of the electron, 
it diverges owing to the zero distance singularity. It appears in both
nonrelativistic and relativistic computations in classical electrodynamics.
We call the zero-distance singularity in quantum local field theories,
relativistic or nonrelativistic,
the ultraviolet infinity. Renormalization theory in perturbative
calculations should be
applied to remove the singularity in a consistent way, respecting
Lorentz and gauge invariance, as well as the quantum principle in the
form of the second quantization.

The difference between non-renormalizable and renormalizable
field theories is only the number of counter terms needed to
remove the ultraviolet infinities; thus, renormalizable theories
have greater predictability in comparison with non-renormalizable
theories, where the number of counter terms rises with the number
of quantum loops. Although the local quantum theory with gravitons
is a non-renormalizable theory, it can be treated with a renormalization
theory just like quantum electrodynamics, but with a necessary 
addition of an increasing number of amplitudes to resolve the large number of
counter terms.

We can conclude that, contrary to the usual wisdom, relativity and
quantum principles are not incompatible in any respect and 
certainly not with respect to the zero-distance singularity
(ultraviolet infinity) in any gauge or non-gauge field theories,
because the quantum principle deals with energies, relativity with
velocities and the ultraviolet singularity is a problem
of the zero-distance between physical entities.
Physical infinities, unlike mathematical ones, usually have 
a physical dimension and a clear physical meaning, as well as
physical consequences. The ultraviolet catastrophe of the
black body radiation and its solution by Planck is an instructive
example.

The second problem is the physical nature of the force of gravity.

The issue of the nature of the force of gravity is closely 
related to the unification programs. The universal law 
of gravity formulated by Newton equals the forces between massive bodies on
Earth with the forces between planets.
Maxwell's theory of electrodynamics unifies the forces of electricity and
magnetism with the theory of light.

The third attempt of unification by Kaluza and Klein, and later by Einstein,
was to unify the forces of gravity and electromagnetism. We now know 
that these early attempts were unsuccessful.
The origin of the electromagnetic force is in an Abelian local gauge force
$U(1)$. The revolutionary consequence of Maxwell's theory is the
existence of the electromagnetic waves, vector-type waves as a
solution to the wave equations with charged sources. The experiments performed
by Hertz proved the existence of waves connected to the electric and
magnetic forces, with characteristics precisely decribed by Maxwell's theory.

By complete analogy with electrodynamics, the theorists assume 
the existence of gravitational waves in Einstein's theory of
gravity. Gravitational waves are solutions to the wave equations.
If the gravitational waves exist, then Einstein's field equations
should have the form of the wave equations. However, from the work of Sciama,
Waylen and Gilman \cite{Sciama1}, it is clear that Einstein's
equations can be put in the form of the generally covariant
integral equations. Thorne put Einstein's equations into the very 
suggestive quasi-wave-type equations, but they evidently remain what
they are, coupled nonlinear integral equations \cite{Thorne}.
There are nonlinear wave equations in electrodynamics because of the
feedback to sources. They are possible because gauge degrees of 
freedom in electrodynamics are internal degrees of freedom, which are completely
distinct from the spacetime degrees of freedom that define 
propagation of waves. This is not the case in gravity, where
internal gauge degrees of freedom coincide with the propagation
spacetime degrees of freedom, making it impossible to form wave equations
from Einstein's field equations. Only empty space (vacuum) and 
pure electromagnetic sources, as unrealizable and unphysical
sources, could lead to the Robinson-Trautman type of gravity wave equations
\cite{Kramer}.

Although one can conclude that no exact wave equation exists for any
realistic physical source, the theorists believe in its existence because,
for example, it is possible to form wave equations for some form
of linearized Einstein's gravity. This is wrong reasoning because
any approximate form of general relativity cannot be a substitute for
the general covariance and relativity principles of the
genuine theory of general relativity. 
One has to prove the existence of the wave equation for the exact form
of GR and then make a certain kind of approximation. The inverse
procedure, when we make the approximation of Einstein's gravity
and then claim that we have a wave equation, is obviously not a proof
of the existence of gravitational waves.

Moreover, I showed in 
ref.\cite{Palle1} that the slowing of the period in the binary pulsar
systems can be understood within the perturbation theory of
general relativity.  The kinetic energy loss of the isolated 
binary pulsar system is compensated by the gain in potential
energy. The interpretation that the kinetic energy loss is compensated
by the gravity radiation is then considered as an indirect proof
of the existence of gravitational waves. These two scenarios differ
obviously at the second order of perturbation theory, but the corrections
are too small to be observed at present.

It is interesting to mention that Cooperstock \cite{Cooperstock}
proved an inability of the gravity waves to carry energy and 
momentum for certain sources. He used a properly defined energy-momentum tensor
and its integrals without referring to any pseudotensor quantities. 
It means that the gravity waves, even for simple and well understood sources, are not
physically observable entities.

From the basic mathematical and physical reasoning, I conclude that gravity is not a local gauge force and that the unification program based
on this assumption will fail. Thus, a quantization of gravity should
not be pursued through the second quantization of the classical
local tensor field, whose existence cannot be established. 

\section{The misuse of the theory of vacuum}

One of the most important concepts in quantum physics is the concept
of vacuum described in Dirac's hole theory \cite{Dirac}.
The vacuum state in QED is defined as a state filled with negative
energy electrons and the Pauli exclusion principle secures the stability
of the ground state. The inevitable consequence is a prediction of
the existence of antiparticles, like positrons in QED. The discovery of
positrons and the success of the QED perturbation theory guarantee
our confidence in Dirac's definition of vacuum.

Faced with new problems, the theorists have to develop new concepts and
mechanisms to solve the obstacles. Let us review a few examples
where new concepts, albeit widely accepted by the physics community, are not physically and mathematically convincing.

The first example is the Higgs mechanism, described by Englert, Brout
and Higgs \cite{Englert,Higgs}. They start with a certain classical scalar
field that has imaginary mass. It could be shown that, at the tree
level for a quantum field and the special choice of the selfinteracting
quartic potential for scalars, the mass of the scalar particle becomes a
real number. This is achieved by redefinition of the scalar field in order to
preserve the vanishing vacuum expectation value of the scalar. Gauge bosons and
fermions acquire masses if they are coupled to this scalar field that
should belong to the scalar representation of the underlying gauge group.
Thus, the scalar field with its interactions plays the role of a
kind of a deus ex machina to solve the mass problem of elementary
particles but at a very high price. Although the mass problem is not
solved because all masses are free parameters, we can at least perform
perturbative calculations even for the spontaneously broken symmetry gauge field
theories.

The second example is very illustrative because it has a few steps of
reasoning before reaching the final physical consequences: (1) BPST
\cite{BPST} found four dimensional Euclidean classical Yang-Mills solutions, i.e. acausal solutions, because we are living in Minkowski spacetime,
(2) 't Hooft calculated a quantum tunneling process based on the BPST
solutions, with the imaginary time resulting in the violation of
chiral symmetry \cite{tHooft}, (3) Peccei and Quin \cite{PQ} assume
that, if the theory contains global $U(1)$ symmetry, the problematic 
P and CP violating terms induced by BPST pseudoparticles
could be removed; and finally, (4) Weinberg \cite{Weinberg1} concludes
that the psudoscalar axion then must exist.
Therefore, starting from the unphysical BPST pseudoparticles living
in the four dimensional Euclidean space, and by the tunneling with
the imaginary time, we are faced with a consequence of the existence
of the pseudoscalar particle axion with unknown mass and couplings. Certain 
models predict that this particle should be the most important
particle in the Universe dominating its present mass density.

The third example represents the radical way how to apparently
resolve the problems of initial conditions in cosmology, i.e.
introduction of the inflationary cosmology.
It is assumed that the dynamics of one or more inflaton scalar fields defines
the fate of the Universe at the GUT scale. This is achieved by the
miracolous adjustement of the selfinteracting scalar potentials,
a game similar to the adjustement of the scalar potentials and
Yukawa couplings in the Higgs mechanism. Although 
the theory of quantum gravity is not established and the treatment
of quantum fields in curved spacetime is far from being an absolved 
subject, the theorists predict the consequences of the inflationary 
cosmology from Planck time to the present. New astrophysical data
are being repeatedly accomodated with new and more complicated scalar potentials
of new inflationary models.

However, all three examples have a lot in common: the problems are
apparently resolved by assuming a transition to a new artificially
constructed vacuum accompanied by a transition from imaginary
to real mass, or imaginary time quantum tunneling of the acausal
instanton solutions, or with an ad hoc treatment of inflaton scalar fields
at Planck and later epochs. Evidently, a clear picture of Dirac's
vacuum, built from quantum and relativity physics, is lost.
 
It would be fair to say, from the historical point of view, that it
is advantageous to have a certain practical and pragmatical mechanism to
gain some insight into the physical realm rather than be completely
helpless. At least we can make calculations in particle
physics and cosmology, but the merit and the background of the above three
examples look more like fabulous adventures of Baron
M\"{u}nchhausen. How to solve the tantalizing problems of fundamental
interactions will be explained in the next chapter, where
the idea of bootstrapping, experienced also by Baron M\"{u}nchhausen,
plays a vital role.

\section{The noncontractible space and the physical world
of finiteness}

Any new theory in particle physics and cosmology has to fulfil
many requirements with respect to the basic physical principles, mathematical
consistency and phenomenological applications.

My fascination with particle physics started when I learned as 
a student about violation of parity in weak interactions.
Why should the world of elementary particles be asymmetric
with respect to the mirror symmetry?

Having learned all about Weinberg-Salam model, 't Hooft-Veltman work etc.,
I still did not understand why the world was asymmetric.
How can the spinless Higgs boson be responsible for violation
of discrete symmetries and the generation of large masses of heavy
quarks and small masses of leptons at the same time?
It reminds me of the phlogiston theory of the pre-Lavoisier chemistry
or of the theory of aether of the pre-Einstein physics.

Could we learn anything useful from the higher dimensional theories or
from the supersymmetric theories?

What could be the guiding principle in an attempt to solve
the problem of elementary particles' masses?

To find an answer, let us in the beginning summarize all our
previous reasoning: 1. gravity is not a local gauge force
and any attempt to unify it with the standard model gauge forces
cannot succeed, 2. elementary scalar fields, such as the Higgs scalar
and the inflaton scalar in particle physics and cosmology, should
be avoided in construction of the theories of fundamental 
interactions, 3. the theory of quantum gravity should be introduced,
4. an alternative mechanism of the gauge symmetry breaking must
be postulated.

I shall now expose my theory as a possible solution to the 
many problems in particle physics and cosmology:

1. The world of particle physics is defined by all possible gauge 
symmetries allowed by conformal $SU(3)$ family as its imbedding symmetries: the whole $SU(3)$ and the $SU(2)\times U(1)$ subsymmetry 
\cite{Palle2}. The number of allowed gauge degrees of freedom in the
conformal space exactly
matches that of the realized strong and electroweak gauge degrees 
of freedom in Minkowski space $6\times 8=4\times 8+4\times 3
+4\times 1$. Therefore, we need the higher dimensional conformal space
only to solve the $SU(2)$ global anomaly problem \cite{Palle2}
and as a scheme for the strong-electroweak unification.
The two additional dimensions are only the dimensions of the auxilliary parameters and 
not real physical dimensions. However, conformal space appears 
naturally as a space of all conformal transformations of 
Minkowski spacetime \cite{Mack}.

2. The SU(2) global anomaly problem can be solved only by the
exact cancellation between the weak gauge boson effective action and
that resulting from fermions (quarks or leptons) interacting
with weak gauge bosons \cite{Palle2}. The negative sign in front
of the second effective action appears because of the functional
integration over fermion fields as grassmannian variables.
A kind of global supersymmetry is present between already
observed elementary fermions and gauge bosons, but not forming new
supermultiplets with particles of 0, 1/2, 1, 3/2 or 2 spin.

3. Mixing angles of weak bosons and fermions (quarks or leptons)
are correlated \cite{Palle2} as a consequence of the invariance of the functional
measure of the electroweak theory on the flavor and boson rotations.

4. Chirally asymmetric couplings of fermions and weak bosons
appear as a necessity of the mathematical consistency of the 
BY theory of \cite{Palle2}, as well as of the Majorana nature
of neutrinos.

5. Instead of the Higgs mechanism, I introduced a noncontractible space
as a symmetry breaking mechanism. It assumes that the spacelike domain
of the four dimensional Minkowski spacetime has lower bound (upper bound
in the Fourier transformed spacetime) fixed at tree level within
the relativistic quantum theory with the non-Abelian quartic selfcouplings
of weak gauge bosons \cite{Palle2}. Thus, the conformal, discrete
and gauge symmetries are broken as a consequence of the assumed
property of the space. This is very appealing because we have a common
comprehension on how to simultaneously break spacetime symmetries (conformal
and discrete) and gauge (internal) symmetries by mass terms generated
in the local relativistic quantum field theory in the noncontractible space.

6. The fermion masses are calculable via bootstrap Dyson-Schwinger equation
(the first instance of the Baron M\"{u}nchhausen bootstrapping). There
is no arbitrariness like with Yukawa couplings. On the contrary, mass functions
of particles
and all the other observables are defined by their gauge invariant 
couplings \cite{Palle2,Palle3}. The appearance of
light and heavy Majorana neutrinos and
other properties of the spectrum is discussed in detail in ref.\cite{Palle2}.

7. The perturbative treatment of the UV finite BY theory \cite{Palle2} of
strong and electroweak interactions is explained
in \cite{Palle4}. The recipe is straightforward and fulfils Lorentz,
gauge and translational symmetries. The UV cutoff is the universal
physical constant that should be extracted by the fit of experimental data by
the formulas of the BY theory.

8. Since gravity is not a local gauge force and Einstein's
gravity is an incomplete theory, I choose Einstein-Cartan theory
of gravity, formulated by Sciama and Kibble \cite{Sciama2}, as
my favorite classical and quantum theory of gravity.

9. The problem with general relativity is that it does not
include rotational degrees of freedom of spacetime and matter.
It is possible to construct the angular momentum in GR but
not as a tensor quantity \cite{Weinberg2}. If one allows that
the linear affine connexion has a symmetric and an antisymmetric part
as an object within Riemann-Cartan geometry, a new relation emerges
between torsion of spacetime and angular momentum including spin
of matter.

10. Torsion is coupled to the total angular momentum; thus,
the quantum mechanical spin of matter also enters the algebraic equations.
The quantum theory acts in Einstein-Cartan gravity on the first
quantized level through spin terms that vanish in the classical
limit. The classical part of the angular momentum is always present, influencing torsion of spacetime even in the classical limit
of the vanishing Planck constant.

11. Spin-torsion effects can avoid the zero-distance singularity
\cite{Trautman,Palle5} in cosmology, with the minimal cosmic scale factor fairly compatible with the universal UV cutoff in particle physics \cite{Palle2}.
If the assumption of the noncontractible space is correct, then
there is no singularity within black holes and spin densities of 
matter configure themselves as a bouncing force to prevent a
collapse beyond the critical universal distance (scale).

12. At this stage, it would be fruitful to comment on the particle
creation process by black holes proposed by Hawking. He claims
\cite{Hawking} that, in the absence of a deeper theory in which
spacetime itself is quantized, one should be satisfied with
an approximation, where the spacetime metric is treated classically
but is coupled to the quantum mechanically treated matter fields.
His starting equation is then the wave equation \cite{Hawking}, namely for scalar fields: $\phi_{;ab}g^{ab}=0$.
This is in analogy with the coupling of the external classical
electromagnetic field with quantum matter fields,
but as previously exposed, this kind of analogy is misleading and wrong.
There is no local quantum gravity tensor field, and the quantum
principle is built in Einstein-Cartan gravity only on the first
quantized level through the matter spin densities. Since there are no
Hawking type wave equations in Einstein-Cartan gravity, Hawking radiation by black holes is forbidden. This is
consistent with our physical insight into the absence of any direct
local physical process between gravity and particles.
The Minkowski spacetime physics and the Riemann-Cartan spacetime physics
are linked only indirectly through Einstein-Cartan equations
and tetrad fields.

13. The cosmological constant problem presents the most elusive
problem in physics, and any attempt to understand
and solve it, requires a reference to the theory of everything,
i.e. to the theory of all fundamental interactions.
The standard wisdom assumes an incredible mixture of concepts,
from the local quantum and classical field theory and GR up to Planck scale, 
various unification schemes, quantum gravity etc., as it is neatly reviewed in
ref.\cite{Weinberg3}. If the cosmological constant is defined as
the zero point energy of a local field or the vacuum energy,
the theoretical calculus
overestimates the observed value by 40, 50 or even more orders of magnitude
\cite{Weinberg3}.
This is a clear signal that a strong departure from the usual wisdom
is necessary. It can be achieved by making a strict distinction between
the local structure of spacetime described by BY theory \cite{Palle2}
(Minkowski spacetime, second quantization of local fields) and the global structure
of spacetime described by Einstein-Cartan gravity \cite{Sciama2}
(Riemann-Cartan spacetime, first quantization) as UV finite field
theories (nonsingular with respect to the zero-distance singularity). 
We can expect that the cosmological constant problem should be
solved within Einstein-Cartan (EC) cosmology. The additional rotational
degrees of freedom of EC gravity provide a kind of bootstrap
(the second instance of Baron M\"{u}nchhausen bootstrapping) at
spacelike infinity to fix the normalization of the mass-density
of the Universe in the model with expansion, acceleration and vorticity
\cite{Palle5}: $\lim_{R\rightarrow \infty}\rho_{m}/\rho_{\Lambda}=-2,
\ \kappa\rho_{m}(R=\infty)=6 H^{2}(R=\infty),\ \kappa=8\pi G_{N}c^{-4}$.
Namely, the number density
of matter particles appears in both EC equations: curvature vs. 
energy-momentum and torsion vs. spin-angular momentum, and the
same coupling constant $\kappa$ figures in both equations.
To conclude, the cosmological constant vanishes:
$\rho_{\Lambda}=-\frac{1}{2}\lim_{R\rightarrow \infty}\rho_{m}=0$.

14. It is important to address the question of how to solve
classic cosmological problems, such as the horizon problem,
the flatness problem and the structure formation problem
within EC gravity and without inflaton scalars.
In the EC gravity, the global structure of spacetime is completely
defined by the matter content of the Universe, including all 
its physical processes, but the reverse statement is also valid:
the expansion and vorticity influence, for example, abundances
of the surviving species of particles, etc.
At the earlier stages of the evolution, spin densities not only
help to avoid the initial singularity \cite{Palle5},
but also trigger the initial primordial density contrast, which
is a necessary ingredient for the structure-formation \cite{Palle6}.
Assuming lepton CP violation, the light neutrino spin densities
induce primordial vorticity and subsequent growth of the angular
momentum of large scale structures \cite{Palle7}, and ultimately
the torsion of spacetime.
The mass-density normalization at
the final evolutionary stage at $T_{\gamma}=0$ \cite{Palle5} tells us that
the limiting ($R\rightarrow \infty \equiv T_{\gamma}\rightarrow 0$)
effective mass-density is the critical one, where the mass-density
is twice the critical one, and the limiting effective contribution of
torsion terms (quadratic and linear) is minus one critical density.
The question of the horizon problem, why a large number of
causally disconnected regions have the same CMB temperature, has a
very clear answer within EC cosmology: local physical processes
in any patch of the Universe are the same and the whole
Universe acts in its patches like a global force. Thus, any
cosmic observable of homogeneity or deviation from homogeneity or
isotropy is global by definition and in this respect well defined.
This is in accord with Machian reasoning \cite{Mach}.

15. The central questions of modern cosmology - what is dark matter
and what is dark energy - in my theory have unique answers:
(1) heavy Majorana neutrinos are cold dark matter particles, because
they are cosmologically abundant and stable $\tau_{N_{i}} \gg 
\tau_{U}$ \cite{Palle8} (the Higgs mechanism generated heavy Majorana
neutrinos cannot be cosmologically stable), (2) angular momentum
(acting like a torsion) of the Universe is dark energy and 
its evolution with redshift and its clustering with dark matter halos \cite{Basilakos,Bean} are to be expected. 

16. In the paper \cite{Palle9}, I argue two possible scenarios
to solve the small CMB power at large scales and large peculiar
velocities of clusters, assuming that the present Universe is
in the vicinity of spacelike infinity $\rho_{\gamma,0}/\rho_{m,0}={\cal O}
(10^{-4})$:
(1) a small Hubble constant and a small contribution
of torsion at low redshifts, and (2) a large Hubble constant and a large 
contribution of torsion at low redshifts. It seems that observations 
favor the large Hubble constant, while the theory requires a large
torsion in the vicinity of infinity \cite{Palle5}. Therefore,
future observations must observe the redshifting of the clustered 
angular momentum of the Universe as dark energy.

17. Now I can go back to the problem of the broken parity in weak
interactions, i.e. the existence of only left handed weak
currents. Why did nature choose the left-handed and not the right-handed
weak currents? The resolution of this dilemma lies in the
insight into the complete local and global structure of spacetime:
the index theorem and homotopy in particle physics require the left-handed
weak currents \cite{Palle2}, while the left-chirality of
weak interactions, together with the violation of the lepton
number and the leptonic and baryonic CP violation, results in the right-handed chirality
of the vorticity of the Universe \cite{Palle7} and the total
chirality vanishes. It means, in other words, that if we chose the
left-handed coordinate system, weak currents would be right-handed and the vorticity of the Universe left-handed. Thus,
particular chiralities are only the substance of our conventions.
Moreover, for the Universe to exist, the broken parities in
particle physics and cosmology must be a mathematical and phenomenological
necessity. The absence of the zero-distance singularity is the
ultimate condition for the existence of the physical world.

\section{Epilogue or facing with the brutal physical reality}

There are great expectations from new experiments and observations
in particle physics, astroparticle physics and cosmology, such
as LHC, neutrino oscillation experiments, $0\nu 2\beta$
decay experiments, AMS at ISS, Auger observatory, LSST, IceCube,
underground detectors for DM-baryon interaction, imaging
atmospheric \v Cerenkov telescopes, Planck and Herschell missions, James Webb
telescope, etc.
At the dawn of probable discoveries and surprises, I have to list 
phenomenological predictions of my theory of the noncontractible
space in more detail:

1. The absence of the asymptotic freedom in QCD embedded in the
noncontractible space $\lim_{\mu \rightarrow \infty} \alpha_{s}^{\Lambda}\neq 0$
means larger QCD amplitudes starting from $\mu \geq 200 GeV$
\cite{Palle4}. The papers referenced in \cite{Palle4} of Tevatron and
especially ref. \cite{Tevatron}, where the quotient of jet cross sections 
at two different center of mass energies free of systematic errors is measured,
strongly suggest a larger QCD coupling than in the Standard Model (SM) at
larger scales. QCD loop corrections to the electroweak processes and a 
deviation due to the larger QCD coupling, are probably marginally observed
by HERA and LEP II \cite{Palle4}. In the year 2000, certain experiments
at LEP II ($\sqrt{s}\simeq 210 GeV$) claimed a discovery of the Higgs
scalar with a mass $M_{H}\simeq 114 GeV$ ($\sqrt{s}\simeq M_{Z}+M_{H}$),
but it was just a nonresonant QCD enhancement above the scale 200 GeV
and its influence on the electroweak couplings.
A similar phenomenon appeared at HERA in 1996-1997.
The LHC should completely resolve the issue and if the universal cutoff exists,
the LHC can measure it very accurately.

2. The small fine structure constant induces small electroweak quantum
corrections, but the combined data of LEP II, SLC and NuTeV show
the discrepancies in $sin^{2}\Theta_{W}$ and $A_{FB}$ from the SM. The
dependence on the Higgs mass is logarithmic and the dependence on the cutoff
of BY theory in ref.\cite{Palle2} is also logarithmic but
with different functional dependences. Even the measurement of the
muon anomalous magnetic moment shows some deviation from the SM. 
The LHC data gives the opportunity to study in detail the quantum
loop structure of the electroweak sector and the symmetry-breaking
mechanism. Any future linear collider will accomplish this task even better. 

3. Although the oscillations of light neutrinos are well
established, the precise masses, mixing angles and particularly the
CP violating phase in lepton sector and the Majorana or Dirac
nature of neutrinos, must be determined by future experiments.

4. The greatest challenge for the astroparticle and particle physics
will be to identify cold dark matter particles, the heavy Majorana
neutrinos with masses from ${\cal O}(10TeV)$ to ${\cal O}(100TeV)$.
It is not excluded that the LHC discovers the pair of the lightest
of the three heavy species. The galactic center H.E.S.S. source
J1745-290 is a perfect source of gamma rays coming from 
annihilation of the CDM particles.
No time-variability of the source spectrum, point-like
topology of the source and the characteristic power spectrum are
almost impossible to interpret by other astrophysical processes.
This H.E.S.S. source, if coming from the CDM annihilation, refers 
to very high masses of the CDM particles and large annihilation 
cross sections; thus, it differs significantly from the standard expectations
of the supersymmetric models. Recent astrophysical results of
the antimatter search (Pamela, ATIC, etc.) and the diffuse photon
background (Fermi-LAT), are still inconclusive for the indirect 
CDM search. New Auger data confirms the GZK cutoff, meaning that very
rare cosmic rays of the highest energies are supressed owing
to the interaction with the CMB. However, there is still a
possibility that subdominant flux of heavy Majorana neutrinos
produced at very large cosmic distances can cause UHE cosmic rays type
events by annihilation with galactic heavy neutrinos \cite{Palle10}. 

5. A great task for cosmology and astrophysics is to measure 
the separate abundances of dark matter and dark energy (angular
momentum of the Universe).
The $\Lambda CDM$ concordance model is already under scrutiny
because of the observed anomalous anisotropic large scales flows,
small large scale power of the CMB, violation of parity (isotropy), etc.

6. The measurement of the vorticity, its chirality and magnitude 
\cite{Palle7,Palle11}, should be exercised by the CMB (corrected
Ellis-Bruni covariant variables must be used \cite{Palle9}),
 by the Faraday rotation of distant radio sources \cite{Birch},
 by examination of spiral galaxies \cite{Longo}, by the
anisotropic anomalous flow of the clusters of galaxies \cite{Kashlinsky},
by the study of galactic correlation functions, by the study of 
the high-redshift objects, etc. If the vorticity does not vanish,
its imprint is everywhere.

7. LISA mission, originally devoted to catching gravity waves, might
measure solar gravity potentials with high accuracy. The presence
of the anomalous cosmic force as a consequence of the cosmic 
acceleration (this is a new independent cosmic parameter like 
expansion or vorticity) can then be easily established
\cite{Palle12}.

8. There are no tensor mode cosmic matter perturbations in my concept of the Universe,
contrary to the inflationary scenario, for two reasons:
1. gravity waves do not exist, and 2. inflationary epoch does not
exist. However, the gravity wave searches must give the ultimate
negative or positive answer, similarly to Michelson-Morley experiment
agenda to study aether, in order to formulate physical laws.
Anyhow, it is more probable to observe the action of quadrupole potentials
\cite{Palle13} on light rays, than to observe quadrupole radiation.

9. The ultraviolet cutoff is explicitly contained in Green's  
functions of my BY theory, treated perturbatively or nonperturbatively
as a quantum relativistic field theory in the noncontractible space.
In ref. \cite{Palle14} I show that the UV cutoff (minimal length)
can be measured in quantum mechanics as a source of spectral
line broadening, using a formalism of the quantum holonomy operators.
The effect is more pronounced for smaller
scale of the quantum mechanical system, such as in the nuclear transitions.

To conclude, the theorists should be more humble with respect to
mathematics and in their theories use particles, vacua and tensors,
instead of pseudoparticles, pseudovacua and pseudotensors. The predictions
of their theories will then be more trustworthy. Otherwise, we 
shall witness an endless search for gravity waves, Higgs bosons,
inflaton scalars, axions, neutralinos, Hawking radiation, instantons, unparticles,
extra dimensions, ...

\end{document}